# Delay-Aware Robust Edge Network Hardening Under Decision-Dependent Uncertainty

### v1

@@ IV-B3 Impact of network size
Figs. 4(a)-4(b) illustrate how the system size affects the optimal solution. As presented in Fig. 4(a), the total cost decreases when the number of ENs increases because the availability of more edge resources enables the platform to efficiently distribute the workload, ensuring that user requests are served by ENs located closer to them. This improved proximity helps to reduce delay and unmet demand penalties, thus enhancing service quality. On the other hand, Fig. 4(b) reveals that, for the same set of ENs, the total cost increases as the number of areas $I$ increases, which translates to a growth in resource demand. This increase in demand generally leads to higher delays and more unmet demand.
@@ IV-B4 Running time comparison between RDDU and e-RDDU
In the simulation, we set the optimality gap to $0.2\%$ and then compute the average running time over $100$ randomly generated problem instances for each problem size. Table III provides a comparison of the computational time taken by RDDU and e-RDDU for various problem sizes and levels of uncertainty. For smaller and medium-sized networks, both algorithms are capable of producing optimal solutions within a reasonable time. However, as the system size grows larger, the advantage of e-RDDU over RDDU becomes more pronounced due to the former’s ability to generate fewer constraints when linearizing bilinear terms. It is worth emphasizing that the underlying problem is a robust planning problem that does not necessitate real-time computation.
@@
TABLE III: Running time comparison
@@ IV-C Performance Comparison
  • NH: There is no hardening at optimality in the deterministic model presented in Section II-B. The platform minimizes only the workload allocation cost in (11).
  • RAND: We consider a randomized link hardening scheme where the platform randomly selects a subset of links for hardening until exhausting the budget $B$. Specifically, the platform has no preferences on which link and which hardening level to be chosen. Thus, it can successively select links for hardening at random until exhausting the budget. Note that there is at most one hardening level chosen for each link. Thus, if a link has been selected at an iteration, it will be removed from the set of available links for hardening in future iterations.
  • SDDU: The platform optimizes the link hardening decision by solving a stochastic model under decision-dependent uncertainty, where the uncertainty link delay is modeled through a set of decision-dependent scenarios. The SDDU model is presented in Appendix -B.
@@
The evaluation and comparison of the four schemes are conducted based on their average actual costs over the generated scenarios. As shown in Figs. 5(a)5(b), the proposed RDDU scheme outperforms the other schemes significantly in terms of the actual total cost. Both $RDDU$ and $SDDU$ take into account the dependence of uncertainty on the decisions when making the planning decisions, giving them an advantage over the other two schemes. However, SDDU’s costs are higher than RDDU’s since SDDU only considers a set of scenarios within the uncertainty set $\mathcal{D}_{2}^{a}$ while $RDDU$ is robust against the whole set. Furthermore, the performance of $SDDU$ depends on the knowledge of the distribution of the historical data. The cost of the $NH$ scheme remains the same as $\Psi$ increases since NH does not implement any hardening, regardless of the value of $\Psi$. Figs. 5(c)5(d) demonstrate that the hardening cost for $NH$ is consistently zero. The $RAND$ scheme, as defined, always tends to exhaust the budget. Additionally, $RDDU$ suggests greater investment in link hardening than $SDDU$.
Figs. 6(a)6(b) compare the four schemes when varying the maximum delay threshold $\Delta$ and the delay penalty parameter $\rho$. It can be observed that for all schemes, the total cost decreases as $\Delta$ increases (i.e., indicating greater relaxation in the delay requirement). On the other hand, as expected, the total cost increases with an increase in the delay penalty $\rho$. Furthermore, the results demonstrate the superior performance of the proposed $RDDU$ scheme compared to the other schemes.
@@ V Related Work
In recent years, significant research efforts have been dedicated to various aspects of EC, with a significant focus on optimal edge service placement and workload allocation. For example, Pasteris et al. [21] studied multi-service placement in EC to maximize the total reward from serving users’ requests. Reference [22] jointly optimizes edge service placement and request routing, while satisfying multi-type resource capacity constraints. In [23], the authors proposed a reliability-aware dynamic service placement model based on the Markov decision process to minimize the total operation costs while maximizing the accepted request rate. The workload allocation problem was formulated as a MILP in [24] to allocate workloads from different areas to available ENs, with the goal of maximizing the request acceptance rate under strict QoS requirements. In [25, 26], the authors proposed a market equilibrium method for allocating heterogeneous edge resources in a fair and efficient manner. Yang et al. [27] investigated the dynamic resource allocation problem among NFV-enabled edge cloud networks and presented a heuristic-based algorithm to minimize the operation cost while satisfying a low latency constraint. [28, 29] investigated the data center operation problem while considering the renewable energy and battery. The studies above mainly focused on deterministic optimization models for EC without considering any system uncertainty. As a result, their performance strongly depends on prediction accuracy, highlighting the need for considering uncertainties in EC.
@@
Optimizing edge networks under uncertainty has gained significant attention in recent research. In [30], the authors employed stochastic programming to minimize edge resource provisioning cost in the face of demand and price uncertainties. Ouyang et al. [31] introduced a bandit-based service placement scheme for joint minimization of the transmission delay and service migration cost in an online setting. In [32], the authors studied a bandit-based dynamic service placement problem in an online and decentralized manner, aiming to maximize total reward while satisfying the delay requirements. Reference [33] proposed a multi-armed bandit approach to online pricing for heterogeneous edge resource allocation. A two-stage adaptive robust model was proposed in [34] to optimize the service placement and sizing decisions for delay-sensitive services while accounting for demand uncertainty. In contrast to these works, our focus is on the uncertainty of link delay, and how to effectively strengthen network links to proactively regulate the level of uncertainty and ensure high service quality. This problem has been overlooked in the existing EC literature.
@@
There is a growing interest in studying reliability and resiliency in networking and cloud/edge computing. In [3, 35], the authors proposed proactive network resource backup solutions to mitigate the impact of random network failures. Johnston et al. [3] presented a robust backup model designed to protect against random link failures, while Ito et al. [36] proposed a different robust model for probabilistic protection to assist the cloud provider in minimizing the required backup capacity for multi-type resources against random failures of physical machines. A backup network design scheme was proposed in [37] to minimize the total required backup capacity with prior knowledge of the probabilistic distribution function of the required backup capacity. Reference [38] investigated the delay-sensitive virtual network function placement and routing problem, considering resiliency, and formulated it as an integer nonlinear program solved by a heuristic algorithm. A resiliency-aware optimal edge service placement framework was proposed in [39] to minimize the total operation cost under demand uncertainty and random node failures. Unlike the aforementioned studies that primarily aimed to provide protection against random failures, our research focuses on the proactive hardening of network links to manage the level of uncertainty and guarantee high service quality in EC. [40] proposed a two-stage decision dependent distributionally robust optimization framework to capture the interdependence between EN placement decisions and uncertain demands. However, this framework also neglects the uncertain network delay as well as dependencies with link utilization.
@@
To the best of our knowledge, we are the first to examine link delay uncertainty in the context of EC and integrate it into a novel robust budget-constrained link hardening model. In contrast to existing robust models that treat uncertainties as exogenous, our model incorporates an endogenous uncertainty set to accurately capture the interdependence among network link delay, link hardening, and workload allocation decisions. By utilizing this endogenous uncertainty set, our model empowers the platform to make decisions that optimize system performance while proactively controlling the level of uncertainty. We also developed efficient algorithms to solve the formulated robust link hardening problem under endogenous delay uncertainty.
@@ VI Conclusion
This paper presented a novel framework for robust link-hardening in EC to mitigate the impact of delay uncertainty on service quality. The main novelty lies in the integration of a DDU set into the proposed framework, which effectively captures the interdependence between uncertain link delay and link hardening as well as workload allocation decisions. Furthermore, we devised two exact algorithms to efficiently compute the optimal solution. Extensive numerical results underscore the advantages of accounting for endogenous uncertainties and demonstrate the superiority of the proposed model over benchmark schemes. The proposed model emphasizes the significance of considering decision-dependent uncertainties when making robust link-hardening decisions.

### v2

@@ IV-B3 Impact of network size
Figs. 6(a)-6(b) illustrate how the system size affects the optimal solution. As presented in Fig. 6(a), given a fixed budget, the total cost decreases when the number of ENs increases because the availability of more edge resources enables the platform to efficiently distribute the workload, ensuring that user requests are served by ENs located closer to them. This improved proximity helps to reduce delay and unmet demand penalties, thus enhancing service quality. In addition, the cost decreases rapidly at first, then more gradually later. This implies that the platform does not require too many ENs to maintain service quality, as unmet demands and delays decrease with the availability of more edge resources. On the other hand, Fig. 6(b) reveals that, for the same set of ENs, the total cost increases as the number of areas $I$ increases, which translates to a growth in resource demand. This increase in demand generally leads to higher delays and more unmet demand. This situation may reduce the platform’s flexibility in resource allocation, potentially leading to unmet demand penalties from delay and resource capacity constraint violations. In contrast, increasing ENs with a fixed number of APs reduces backhaul pressure and provides more flexibility in allocating services to nodes, lowering total cost.
@@ IV-B4 Running time comparison between RDDU and e-RDDU
In the simulation, we set the optimality gap to $0.2\%$ and then compute the average running time over $100$ randomly generated problem instances for each problem size. Table IV provides a comparison of the computational time taken by RDDU and e-RDDU for various problem sizes and levels of uncertainty. For smaller and medium-sized networks, both algorithms are capable of producing optimal solutions within a reasonable time. However, as the system size grows larger, the advantage of e-RDDU over RDDU becomes more pronounced due to the former’s ability to generate fewer constraints when linearizing bilinear terms. It is worth emphasizing that the underlying problem is a robust planning problem that does not necessitate real-time computation.
@@
TABLE IV: Runtime comparison
@@ IV-C Performance Comparison
  • $\bullet$ NH: There is no hardening at optimality in the deterministic model presented in Section II-B. The platform minimizes only the workload allocation cost in (2).
  • $\bullet$ RAND: We consider a randomized link hardening scheme where the platform randomly selects a subset of links for hardening until exhausting the budget $B$. Specifically, the platform has no preferences on which link and which hardening level to be chosen. Thus, it can successively select links for hardening at random until exhausting the budget. Note that there is at most one hardening level chosen for each link. Thus, if a link has been selected at an iteration, it will be removed from the set of available links for hardening in future iterations.
  • $\bullet$ SDDU: The platform optimizes the link hardening decision by solving a stochastic model under DDU, where the uncertainty link delay is modeled through a set of decision-dependent scenarios. The SDDU model is presented in Appendix -B.
@@
The evaluation and comparison of the four schemes are conducted based on their average actual costs over the generated scenarios. As shown in Figs. 7(a)7(b), the proposed RDDU scheme outperforms the other schemes significantly in terms of the actual total cost. Both $RDDU$ and $SDDU$ take into account the dependence of uncertainty on the decisions when making the planning decisions, giving them an advantage over the other two schemes. However, SDDU’s costs are higher than RDDU’s since SDDU only considers a set of scenarios within the uncertainty set $\mathcal{D}_{2}^{a}$ while $RDDU$ is robust against the whole set. Furthermore, the performance of $SDDU$ depends on the knowledge of the distribution of the historical data. The cost of the $NH$ scheme remains the same as $\Psi$ increases since NH does not implement any hardening, regardless of the value of $\Psi$. Figs. 7(c)7(d) demonstrate that the hardening cost for $NH$ is consistently zero. The $RAND$ scheme, as defined, always tends to exhaust the budget. Additionally, $RDDU$ suggests greater investment in link hardening than $SDDU$.
Figs. 8(a)8(b) compare the four schemes when varying the maximum delay threshold $\Delta$ and the delay penalty parameter $\rho$. It can be observed that for all schemes, the total cost decreases as $\Delta$ increases (i.e., indicating greater relaxation in the delay requirement). On the other hand, as expected, the total cost increases with an increase in the delay penalty $\rho$. Furthermore, the results demonstrate the superior performance of the proposed $RDDU$ scheme compared to the other schemes.
@@ V Related Work
This section provides an overview of existing research in edge computing, focusing on service placement, workload allocation under uncertainty, and network reliability, highlighting how our work extends and differs from previous studies.
@@
Edge service placement and workload allocation: Extensive research has recently focused on optimizing edge service placement and workload allocation. Pasteris et al. [24] studied multi-service placement in EC to maximize rewards from serving user requests. Reference [25] jointly optimized service placement and request routing under resource capacity constraints. A reliability-aware dynamic service placement model based on the Markov decision process was proposed in [26] to minimize operational costs while maximizing accepted requests. The workload allocation problem was formulated as a MILP in [27], aiming to maximize request acceptance while meeting strict QoS requirements. References [28, 29] proposed market equilibrium methods for fair and efficient edge resource allocation, and [30] explored the data center operations considering renewable energy and batteries. However, most of these studies rely on deterministic optimization models, making their performance heavily dependent on prediction accuracy, highlighting the need to incorporate uncertainty in EC systems.
@@
Edge resource management under uncertainties: Optimizing edge networks under uncertainty has garnered considerable attention in recent research. Reference [31] used stochastic programming to minimize edge resource provisioning costs amid demand and price uncertainties. Ouyang et al. [32] introduced a bandit-based service placement scheme to jointly minimize transmission delay and service migration costs in an online setting. In [33], the authors studied a bandit-based dynamic service placement problem in a decentralized manner, aiming to maximize total reward while satisfying delay requirements. [34] proposed a multi-armed bandit approach for online pricing in heterogeneous edge resource allocation. A two-stage adaptive robust model was proposed in [35] to optimize edge service placement and sizing decisions under demand uncertainty. Unlike these works, our focus is on link delay uncertainty and how to effectively strengthen network links to proactively manage this uncertainty and ensure high service quality. This problem has been largely overlooked in existing EC literature.
@@
Edge reliability and resiliency: There is growing interest in studying reliability and resiliency in networking and cloud/edge computing. In [3, 36], the authors proposed proactive network resource backup solutions to mitigate the impact of random network failures. Ito et al. [37] introduced a probabilistic protection model to minimize the backup capacity required by cloud providers in the event of random physical machine failures. In [38], a backup network design scheme was presented to minimize the total required backup capacity with prior knowledge of the probabilistic distribution of required capacity. A resiliency-aware optimal edge service placement framework was proposed in [39] to minimize operational costs under demand uncertainty and random node failures. Unlike these studies that primarily focus on protection against random failures, our research emphasizes proactive network link hardening to manage uncertainty and guarantee high service quality in EC. While [40] proposed a decision-dependent distributionally RO model to capture the interdependence between EN placement decisions and uncertain demands, it overlooks uncertain network delays and their dependencies on link utilization.
@@
To the best of our knowledge, we are the first to examine link delay uncertainty in EC and incorporate it into a novel robust budget-constrained link hardening model. In contrast to existing robust models that treat uncertainties as exogenous, our approach introduces an endogenous uncertainty set to accurately capture the interdependence between network link delay, link hardening, and workload allocation decisions. By leveraging this endogenous uncertainty set, our model empowers the platform to make decisions that optimize system performance while proactively controlling uncertainty levels. We also developed efficient algorithms for the formulated robust link hardening problem under endogenous delay uncertainty.
@@ VI Conclusion
This paper introduced a novel framework for robust link-hardening in EC to mitigate the impact of delay uncertainty on service quality. The main novelty lies in the integration of a DDU set, which effectively captures the interdependence between uncertain link delays and link hardening and workload allocation decisions. Furthermore, we devised two exact algorithms, RDDU and e-RDDU, to efficiently compute optimal solutions. The enhanced robust solution, e-RDDU, significantly improves computational speed by eliminating redundant constraints without sacrificing solution optimality. Extensive numerical results underscore the advantages of accounting for endogenous uncertainties and demonstrate the superiority of the proposed model over benchmark schemes. The proposed model emphasizes the significance of considering DDU when making robust link-hardening decisions. In future work, we aim to leverage data-driven approaches, such as machine learning and deep learning techniques, to design fast heuristic algorithms that accelerate the computation of robust solutions for the proposed model. Our current methods, which generate exact solutions, can provide high-quality datasets for training, thereby enhancing predictive accuracy.